OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line. (0,5); 7x+2y=9

OpenStudy (anonymous):

let 2x -7y=k put (0,5) k=-35 2x-7y+35=0

OpenStudy (anonymous):

that doesn't make sense. shouldn't it be 7x+2y??? not 2x-7y?

OpenStudy (anonymous):

its the rule..that product of slope of two perpendicular line is -1

OpenStudy (anonymous):

let x = 0 and y = 5 so, 7(0)+ 2(5) = 9 0 + 10 = 9 then product of slopes -1 + 10 = 9 9=9

OpenStudy (anonymous):

the first line has slope -7/2 and the perpendicular line must have slope 2/7 thats why i written 2x-7y=k

OpenStudy (anonymous):

but the equation of the line needs to be in y= form. what would that be?

OpenStudy (anonymous):

2x-7y+35=0 7y=2x+35 y=2/7x+5